OpenStudy (anonymous):

Evaluate. \[\lim_{x \rightarrow -\infty} \left[ \frac{ 2x }{ \sqrt{4x^2+32x}+x } \right]\]

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}\lim_{x \rightarrow -\infty} \left[ \frac{ 2x }{ \sqrt{4x^2+32x}+x } \right] \hspace{.33em}\\~\\ \implies \lim_{x \rightarrow -\infty} \left[ \frac{ 2 }{ \sqrt{4+\dfrac{32}{x}}+1 } \right] \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (zzr0ck3r):

\(\lim_{x \rightarrow -\infty} \left[ \frac{ 2x }{ \sqrt{4x^2+32x}+x } \right]=\lim_{x \rightarrow -\infty} \left[ \frac{ \frac{1}{x}2x }{\frac{1}{x}( \sqrt{4x^2+32x}+x )} \right]=\lim_{x \rightarrow -\infty} \left[ \frac{ 2 }{ \frac{1}{x}\sqrt{4x^2+32x}+1 } \right]\\=\lim_{x \rightarrow -\infty} \left[ \frac{ 2 }{ \sqrt{\frac{4x^2}{x^2}+\frac{32}{x}}+1 } \right]=\lim_{x \rightarrow -\infty} \left[ \frac{ 2 }{ \sqrt{4+\frac{32}{x}}+1 } \right]\) Now run the limit

OpenStudy (anonymous):

okay...

geerky42 (geerky42):

So do you understand what is going on?

OpenStudy (anonymous):

I think so... the answer is 2/3? @geerky42

geerky42 (geerky42):

Yeah

OpenStudy (anonymous):

Thanks!